Adam can paint a house in 5 hours. Audrey can paint the same house in 6 hours. If they work together how long will it take
to paint the house?

Answer by nerdybill(7384)

Adam can paint a house in 5 hours. Audrey can paint the same house in 6 hours. If they work together how long will it take
to paint the house?
.
Let x = time (hours) for them working together
then
x(1/5 + 1/6) = 1
multiplying both sides by 30:
x(6 + 5) = 30
11x = 30
x = 30/11
x = 2.73 hours
or
x = 2 hours and 44 minutes